### v1

@@ Appendix B Deferred Proofs for Section 4
Substituting $N=\left\lceil T^{2/3}(\ln T)^{1/3}\right\rceil$ yields the stated regret bound. ∎
Appendix C Deferred Proofs for Section 5
See 7
Proof of Lemma 7.
Assume towards contradiction that there exists a matrix $P\in\mathcal{C}$ and a pair $(u,k)$ such that $P_{u,k}>P^{t}_{u,k}$. By the suggested explicit form of $P^{t}$, there exists an index $k^{\prime}$ with either $k\preceq k^{\prime}\preceq p(u)$ or $p(u)\preceq k^{\prime}\preceq k$ such that
$P^{t}_{u,k}=UCB_{u,k^{\prime}}(t).$
From our assumption, we have $P_{u,k}>UCB_{u,k^{\prime}}$. Since $P\in\mathcal{C}$, Definition 4 ensures that $P_{u,k^{\prime}}\leq UCB_{u,k^{\prime}}(t)$. Combining these inequalities yields
$P_{u,k}>UCB_{u,k^{\prime}}(t)\geq P_{u,k^{\prime}}.$
However, the single-peaked property of $P$ with peak at $p(u)$ implies $P_{u,k}\leq P_{u,k^{\prime}}$ (since both arms lie on the same side of the peak), contradicting $P_{u,k}>P_{u,k^{\prime}}$, and the proof is complete. ∎
See 8
Proof of Theorem 8.
We condition the analysis on the event that for all rounds $t\in[T]$ and pairs $(u,k)$ it holds that
$LCB_{u,k}(t)\leq\theta_{u,k}\leq UCB_{u,k}(t).$
Due to Hoeffding’s inequality and a union bound over all pairs $(u,k)$ and rounds $t$, this event holds with probability at least $1-\frac{2KU}{T^{3}}$. As the complementary event contributes at most $TU$ to the regret, its effect on the regret is negligible.
Conditioning on this event and summing Inequality (5) over all rounds $t\in[T]$, we obtain
@@ Appendix C Deferred Proofs for Section 5
$R_{T}$ $\leq O\left(\sum_{t=1}^{T}\sum_{u\in[U]}2\sqrt{\frac{2\ln T}{n_{u,\pi_{t}(u)}(t)}}\right)$
@@
$=O\left(\sqrt{\ln T}\sum_{u\in[U]}\sum_{k\in[K]}\sum_{t=1}^{n_{u,k}(T)}\frac{1}{\sqrt{t}}\right)$
$=O\left(\sqrt{\ln T}\sum_{u\in[U]}\sum_{k\in[K]}\sqrt{n_{u,k}(T)}\right).$
Using Jensen’s inequality, we get:
$R_{T}\leq O\left(\sqrt{\ln T}\sum_{u\in[U]}\sqrt{K}\sqrt{\sum_{k\in[K]}n_{u,k}(T)}\right)=O\left(U\sqrt{TK\ln T}\right),$
which completes the regret analysis.
For the runtime analysis, observe that each round of Algorithm 3 requires three main computational steps: constructing the maximal matrix $P^{t}$ according to Lemma 7, computing the optimal matching via SP-Matching, and updating the empirical statistics. The maximal matrix construction requires $O(UK)$ time to compute all entries. By Theorem 2, the SP-Matching algorithm runs in $O(K^{2}U+K^{2}B)$ time. Updating the empirical means and confidence bounds takes $O(U)$ time. Therefore, the per-round runtime is dominated by the SP-Matching call, yielding $O(K^{2}U+K^{2}B)$ time per round. ∎
Remark 3.
@@
Consider the case where some user $u$ has multiple peak indices, and this fact is known to the algorithm beforehand. In this case, the maximal matrix that is consistent with this new information should allocate the same value to all peak indices of user $u$. Additionally, this value is the minimum of the upper confidence bounds of all peak indices. More concretely, given that user $u$ has peak indices $p,p+1,\ldots,p+m$ for some $m\geq 0$, we should set
$P^{t}_{u,k}=\begin{cases}\min_{i\in[0,m]}UCB_{u,p+i}(t),&k\in\{p,p+1,\ldots,p+m\}\\ \min_{i:k\preceq i\preceq p+m}UCB_{u,i}(t),&k\prec p\\ \min_{i:p\preceq i\preceq k}UCB_{u,i}(t),&p+m\prec k\end{cases}.$

### v2

@@ Appendix E Deferred Proofs from Section˜6
Substituting $N=\left\lceil T^{2/3}(\ln T)^{1/3}\right\rceil$ yields $R_{T}\leq O(UKT^{2/3}(\ln T)^{1/3})=\tilde{O}(UKT^{2/3})$, which completes the proof. ∎
See 12
Proof of Theorem˜12.
We reduce from MAX Betweenness [31, 3], which involves, given a set of elements $S$ and a collection of ordered triplets $G\subseteq S^{3}$, finding a linear order of $S$ that maximizes the number of triplets $(a,b,c)\in G$ for which element $b$ lies between $a$ and $c$ in the order. Austrin et al. [3] show that for any $\eta>0$, it is NP-hard to distinguish between instances where at least $(1-\eta)$ fraction of triplets can be satisfied versus at most $(1/2+\eta)$ fraction.
Reduction.
Given a Betweenness instance $(S,G)$, construct a Max-SP-WCS instance as follows. Create an arm for each element $s\in S$, so $K=S$. For each triplet $g=(a_{g},b_{g},c_{g})\in G$, create two users $u_{g}$ and $v_{g}$, so $U=\{u_{g},v_{g}:g\in G\}$. Define the matching by $\pi(u_{g})=a_{g}$ and $\pi(v_{g})=c_{g}$. Fix $\varepsilon\in(0,1/2)$ and set confidence intervals as:
$\textnormal{LCB}_{u_{g},k},\textnormal{UCB}_{u_{g},k}]=\begin{cases}[1-\varepsilon,1],&k=b_{g}\\ [0,\varepsilon],&k=c_{g}\\ [0,1],&\text{otherwise}\end{cases},$
$\textnormal{LCB}_{v_{g},k},\textnormal{UCB}_{v_{g},k}]=\begin{cases}[1-\varepsilon,1],&k=b_{g}\\ [0,\varepsilon],&k=a_{g}\\ [0,1],&\text{otherwise}\end{cases}.$
This construction is polynomial in the size of the Betweenness instance.
Key observation.
Fix any arm order $\prec$ and triplet $g=(a_{g},b_{g},c_{g})$. If $b_{g}$ lies between $a_{g}$ and $c_{g}$ in $\prec$, we can complete both rows $u_{g}$ and $v_{g}$ unimodally with peaks at $b_{g}$ that respect the confidence intervals and achieve $P_{u_{g},a_{g}}=P_{v_{g},c_{g}}=1$: set $P_{u_{g},k}=1$ for all arms $k$ on the same side of $b_{g}$ as $a_{g}$ (including $b_{g}$), and $P_{u_{g},k}=0$ otherwise; similarly for $v_{g}$.
Conversely, if $b_{g}$ does not lie between $a_{g}$ and $c_{g}$, then either $c_{g}$ lies between $a_{g}$ and $b_{g}$, or $a_{g}$ lies between $b_{g}$ and $c_{g}$. In the former case, single-peakedness requires $P_{u_{g},a_{g}}\leq P_{u_{g},c_{g}}$ or $P_{u_{g},b_{g}}\leq P_{u_{g},c_{g}}$; since the confidence intervals enforce $P_{u_{g},c_{g}}\leq\varepsilon$ and $P_{u_{g},b_{g}}\geq 1-\varepsilon$, we must have $P_{u_{g},a_{g}}\leq\varepsilon$. By symmetry, in the latter case, $P_{v_{g},c_{g}}\leq\varepsilon$. These bounds can be achieved with equality.
Therefore, the contribution from users $u_{g},v_{g}$ is exactly $2$ if triplet $t$ is satisfied by $\prec$, and $1+\varepsilon$ otherwise. Hence, if $s$ triplets are satisfied:
$V(\pi;P)=2s+(|G|-s)(1+\varepsilon)=|G|(1+\varepsilon)+s(1-\varepsilon).$
Hardness of approximation.
[…]
$\textsc{OPT}(\hat{I}_{H})\geq|G|(1+\varepsilon)+|G|(1-\eta)(1-\varepsilon)=|G|(2-\eta+\varepsilon\eta).$
On the other hand, if at most $(1/2+\eta)|G|$ triplets can be satisfied, which we refer to as case $L$:
$\textsc{OPT}(\hat{I}_{L})\leq|G|(1+\varepsilon)+|G|(1/2+\eta)(1-\varepsilon)=|G|\left(\nicefrac{{3}}{{2}}+\eta+\nicefrac{{\varepsilon}}{{2}}-\varepsilon\eta\right).$
[…]
$\alpha>\frac{\textsc{OPT}(\hat{I}_{L})}{\textsc{OPT}(\hat{I}_{H})}\leq\frac{\frac{3}{2}+\eta+\frac{\varepsilon}{2}-\varepsilon\eta}{2-\eta+\varepsilon\eta}\xrightarrow{\eta,\varepsilon\to 0}\frac{3/2}{2}=\frac{3}{4}.$
Thus, for any $\delta>0$, a $(\frac{3}{4}+\delta)$-approximation algorithm would distinguish between hard instances for sufficiently small $\eta,\varepsilon$, contradicting the NP-hardness of MAX Betweenness. ∎
Appendix F Deferred Algorithms and Proofs from Section 7
F.1 Details and Proof for Section˜7.1
In this appendix, we provide the full specification and regret analysis for the Sep-MvM algorithm discussed in Section 7.1. We first formally present the algorithm in Algorithm˜5 and then prove Proposition˜13.
Algorithm 5 Sep-MvM
1:Initialize $n_{u,k}\leftarrow 0$, $\bar{\theta}_{u,k}\leftarrow 0$ for all $u,k$
2:// Phase 1: Exploration with Monitoring
3:for $t=1,\dots,\tau=KT^{2/3}$ do
4:  Select arm $k_{t}=(t\pmod{K})+1$ for all users (Round-Robin)
5:  Observe rewards, update empirical means $\bar{\theta}_{u,k}$ and counts $n_{u,k}$
6:  Update Confidence Intervals: $\mathcal{I}_{u,k}(t)=[\bar{\theta}_{u,k}-\sqrt{\frac{2\ln T}{n_{u,k}}},\bar{\theta}_{u,k}+\sqrt{\frac{2\ln T}{n_{u,k}}}]$
7:  $\texttt{SEP}\leftarrow\forall u,\forall i\neq j,\mathcal{I}_{u,i}(t)\cap\mathcal{I}_{u,j}(t)=\emptyset$
8:  if SEP is True then
9:   Break Phase 1., proceed to Phase 2
10:// Phase 2: Structure Exploitation
11:if SEP is True (Structure Recovered) then
[…]
13:  Run MVM using recovered order $\prec$ and inferred peaks
14:else
15:// Timeout Reached - fallback to EMC Strategy
[…]
17:  Commit to $\bar{\pi}$ for rounds $t,\dots,T$
See 13
Proof of Proposition˜13.
We analyze the regret conditioned on the standard clean event $\mathcal{E}=\{\forall u,k,t:|\hat{\theta}_{u,k}(t)-\theta_{u,k}|\leq\sqrt{\frac{2\ln T}{n_{u,k}(t)}}\}$, which holds with probability at least $1-2UKT^{-2}$. Under $\mathcal{E}$, the true mean is always contained within the confidence intervals $\mathcal{I}_{u,k}(t)$.
We consider two cases based on the magnitude of the instance gap $\Delta_{\textnormal{global}}$.
Case 1: $\Delta_{\textnormal{global}}<\sqrt{\frac{32\ln T}{T^{2/3}}}$.
First, notice that if separation does occur before the timeout $\tau=KT^{2/3}$, then the regret will be at most $O(UKT^{2/3})$, as it comprises of the exploration regret (which will be less that $KT^{2/3}$ rounds) and the regret of the MvM algorithm, which is at most $O(U\sqrt{TK})$ when run with the correct parameters. Next, suppose the separation condition is not met by the timeout. The algorithm switches to the EMC commitment strategy, after performing its exact exploration phase with the same exploration parameter as in Theorem˜11. Thus, the exact same regret bound applies and we get $R(T)\leq\tilde{O}(UKT^{2/3})$. And indeed, when the condition $\Delta_{\textnormal{global}}<\sqrt{\frac{32\ln T}{T^{2/3}}}$ holds, this is the smallest of the two expressions in the minimum (up to poly-logarithmic factors).
Case 2: $\Delta_{\textnormal{global}}\geq\sqrt{\frac{32\ln T}{T^{2/3}}}$.
In this case, separation occurs before the timeout. To see this, notice that under the clean event, separation must occur if $\sqrt{\frac{2\ln T}{n_{u,k}}}\leq\Delta_{\textnormal{global}}/4$. Since samples are collected round-robin, $n_{u,k}(t)\approx t/K$. The condition requires:
$\sqrt{\frac{2K\ln T}{t}}<\frac{\Delta_{\textnormal{global}}}{4}\implies t>\frac{32K\ln T}{\Delta_{\textnormal{global}}^{2}}.$
Thus, the separation event occurs at $t_{sep}=O(\frac{K\ln T}{\Delta_{\textnormal{global}}^{2}})<\tau$.
After the separation, the structure can be recovered perfectly (again, conditioning on the clean event)–pairwise disjointness implies the sorted order of empirical means is identical to the sorted order of true means. Thus, Extract-Order will recover a true SP order and peaks.
After $t_{sep}$, we switch to MVM. The total regret is the sum of the exploration regret (up to $t_{sep}$) and the MVM regret (from $t_{sep}$ to $T$):
$R(T)\leq\underbrace{U\cdot t_{sep}}_{\text{Exploration}}+\underbrace{R_{\text{MVM}}(T-t_{sep})}_{\text{MVM Regret}}$
Substituting $t_{sep}=\tilde{O}(K/\Delta_{\textnormal{global}}^{2})$ and the MVM bound $\tilde{O}(U\sqrt{TK})$:
@@ Appendix F Deferred Algorithms and Proofs from Section 7
$R(T)\leq\tilde{O}\left(\frac{UK}{\Delta_{\textnormal{global}}^{2}}+U\sqrt{TK}\right).$
Conclusion.
Taking the minimum of the two cases (since the algorithm automatically executes the better strategy via the timeout mechanism), we obtain the bound:
@@ Conclusion.
$\tilde{O}\left(U\min\left\{KT^{2/3},\sqrt{TK}+\frac{K}{\Delta_{\textnormal{global}}^{2}}\right\}\right).$
F.2 Details and Proof for Section˜7.2
Algorithm 6 NSP-EMC
1:Exploration rounds NN
2:Match each user $u$ to each arm $k\in[K]$ for NN rounds; let $\overline{\Theta}$ be the empirical mean matrix
3:Binary Search for $\varepsilon$:
4:Set $\varepsilon_{low}\leftarrow 0,\varepsilon_{high}\leftarrow 1,\hat{\varepsilon}\leftarrow 1,\prec\leftarrow\text{Identity}$
5:while $\varepsilon_{high}-\varepsilon_{low}>\frac{1}{T}$ do
6:  $\varepsilon_{mid}\leftarrow(\varepsilon_{low}+\varepsilon_{high})/2$
7:  $\prec_{temp}\leftarrow\textsc{Extract-Order}(\overline{\Theta},\varepsilon_{mid})$
8:  if $\prec_{temp}\neq\text{fail}$ then
9:   $\hat{\varepsilon}\leftarrow\varepsilon_{mid}$
10:   $\prec\leftarrow\prec_{temp}$
11:   $\varepsilon_{high}\leftarrow\varepsilon_{mid}$
12:  else
13:   $\varepsilon_{low}\leftarrow\varepsilon_{mid}$
14:Construct SP matrix $\tilde{\Theta}$ from $(\overline{\Theta},\prec)$ via Lemma˜9 with $\delta=2K\varepsilon_{high}$
[…]
16:Play $\tilde{\pi}$ for the remaining TNKT-NK rounds
We formally present the NSP-EMC algorithm (Algorithm 6), which adapts the explore-then-commit framework to non-single-peaked instances. The algorithm proceeds in three main phases. First, it collects empirical reward estimates through uniform exploration. Second, the algorithm performs a binary search to identify the minimal $\varepsilon$ required to recover a valid permutation from Extract-Order. Finally, it constructs a feasible SP matrix based on this minimal $\varepsilon$ and commits to the corresponding optimal matching for the duration of the horizon.
See 14
Proof of Proposition˜14.
Let $P^{\star}\in\mathcal{S}$ be an SP matrix such that $\|\Theta-P^{\star}\|_{\infty}=\gamma$. We again use the “clean event" technique and condition on it for the rest of the analysis.
Feasibility of the Binary Search.
By the triangle inequality and the definition of $\gamma$:
$\|\overline{\Theta}-P^{\star}\|_{\infty}\leq\|\overline{\Theta}-\Theta\|_{\infty}+\|\Theta-P^{\star}\|_{\infty}\leq\Delta_{est}+\gamma.$
Let $\varepsilon^{\star}=\Delta_{est}+\gamma$. Since $P^{\star}$ is an SP matrix (and thus valid with respect to some SP order $\prec_{P^{\star}}$), Lemma 5 guarantees that the call $\textsc{Extract-Order}(\overline{\Theta},\varepsilon^{\star})$ will succeed and return a valid order (or specifically, that the constraints imposed by $\varepsilon^{\star}$ are consistent with $\prec_{P^{\star}}$). Because the binary search in Algorithm 6 finds the minimal feasible $\varepsilon$ (up to precision $\eta$) for which Extract-Order returns a valid order (note that this condition is indeed monotone in $\varepsilon$), the found parameter $\hat{\varepsilon}$ satisfies:
$\varepsilon_{high}\leq\varepsilon^{\star}+\frac{2}{T}=\Delta_{est}+\gamma+\frac{2}{T}.$
We will omit the $\frac{2}{T}$ term for the rest of the analysis, as it is negligible in the final bound.
Bounding the Approximation Error.
The algorithm produces an order $\prec$ using $\varepsilon_{high}$. According to Lemma 5, the empirical matrix $\overline{\Theta}$ is $(2K\varepsilon_{high})$-ASP with respect to $\prec$. Subsequently, the algorithm constructs $\tilde{\Theta}$ using Lemma˜9 with $\delta=2K\varepsilon_{high}$. Lemma˜9 guarantees that $\tilde{\Theta}$ is SP with respect to $\prec$ and satisfies:
$\|\tilde{\Theta}-\overline{\Theta}\|_{\infty}\leq 2K\varepsilon_{high}.$
Substituting the bound for $\varepsilon_{high}$ from Step 1:
$\|\tilde{\Theta}-\overline{\Theta}\|_{\infty}\leq 2K(\Delta_{est}+\gamma).$
Total Value Gap and Regret.
We now bound the difference between the true optimal matching $\pi^{\star}$ and our committed matching $\tilde{\pi}$. Recall that $\tilde{\pi}$ is optimal for $\tilde{\Theta}$.
$V(\pi^{\star};\Theta)-V(\tilde{\pi};\Theta)$ $=V(\pi^{\star};\Theta)-V(\pi^{\star};\tilde{\Theta})+V(\pi^{\star};\tilde{\Theta})-V(\tilde{\pi};\Theta)$
$\leq V(\pi^{\star};\Theta)-V(\pi^{\star};\tilde{\Theta})+V(\tilde{\pi};\tilde{\Theta})-V(\tilde{\pi};\Theta)$
$\leq 2U\|\Theta-\tilde{\Theta}\|_{\infty}$.
Using the triangle inequality on the norm:
$\|\Theta-\tilde{\Theta}\|_{\infty}$ $\leq\|\Theta-\overline{\Theta}\|_{\infty}+\|\overline{\Theta}-\tilde{\Theta}\|_{\infty}$
$\leq\Delta_{est}+2K(\Delta_{est}+\gamma)$
$=(2K+1)\Delta_{est}+2K\gamma.$
Thus, the per-round regret during the commitment phase is bounded by:
$V(\pi^{\star};\Theta)-V(\tilde{\pi};\Theta)\leq 2U(2K+1)\Delta_{est}+4UK\gamma.$
By substituting this bound into the regret and plugging in our choice of $N$, we get:
$R_{T}$ $\leq NKU+T\left(2U(2K+1)\sqrt{\frac{2\ln T}{N}}+4UK\gamma\right)$
$\leq\tilde{O}(UKT^{2/3}+\gamma UKT).$
Appendix G Handling Ties in Peak Indices
In the classical study of single-peaked preferences in social choice theory, preferences are ordinal—each user provides a ranking over alternatives rather than cardinal utility values. In this ordinal setting, ties at the peak do not arise: the peak is simply the top-ranked alternative, which is unique by definition. However, in our cardinal framework, a user $u$ may have multiple indices $k$ that achieve the same maximal expected reward $\max_{k^{\prime}\in[K]}\theta_{u,k^{\prime}}$. When this occurs, the peak $p(u)$ as defined in Definition˜1 is not unique, and we must verify that our algorithms and analyses are robust to arbitrary tie-breaking choices.
Robustness of the Offline Algorithm.
We begin by establishing that Lemma˜4 is robust to the arbitrary choice of peak when a user has several maximizers. Suppose user $u$ has two peak indices $x<y$ with $\theta_{u,x}=\theta_{u,y}=\max_{k\in[K]}\theta_{u,k}$, and fix a selected set $S=\{k_{1}<k_{2}<\cdots<k_{m}\}$. We claim that the value of the optimal matching of $u$ against $S$ is independent of whether we set $p(u)=x$ or $p(u)=y$. To see this, observe that exactly one of the following cases holds:
  • There exists $j\in\{1,\ldots,m-1\}$ with $k_{j}\leq x\leq y\leq k_{j+1}$ and no $k_{\ell}\in(x,y)$. By unimodality, any arm outside $\{k_{j},k_{j+1}\}$ is dominated by one of them, so $\pi^{\star}(u)\in\{k_{j},k_{j+1}\}$ regardless of the tie-break.
  • $y\leq k_{1}$. Values are non-increasing to the right of the peaks, hence $\pi^{\star}(u)=k_{1}$ regardless of the tie-break.
  • $k_{m}\leq x$. Values are non-decreasing to the left of the peaks, hence $\pi^{\star}(u)=k_{m}$ regardless of the tie-break.
  • There exist indices $i\leq j$ with $x\leq k_{i}\leq\cdots\leq k_{j}\leq y$. Since all $k\in[x,y]$ attain the peak value by the single-peaked property, any $k\in\{k_{i},\ldots,k_{j}\}$ is optimal. The lemma asserts $\pi^{\star}(u)\in\{k_{i-1},k_{i}\}$ if the peak is set to $x$, and $\pi^{\star}(u)\in\{k_{j},k_{j+1}\}$ if the peak is set to $y$. In either case, the attained value $\max_{k\in S}\theta_{u,k}$ remains unchanged.
In every case, user $u$’s contribution to the matching value is invariant to the choice of peak.
This invariance extends to SP-Matching (Algorithm˜4). Observe that every choice of the rightmost arm $k$ in Line 9 corresponds to a subset of selected arms and a matching that assigns each user according to Lemma˜4. By the argument above, the attained value of each user in this subset is independent of how we resolve peak ties. The choice of peak only affects when we account for a user’s contribution during the dynamic programming computation, not the contribution itself. Therefore, running SP-Matching with different valid peak indices yields the same optimal value $V^{\star}$.
Robustness of the Online Algorithm with Known Structure.
@@ Appendix G Handling Ties in Peak Indices
For the MvM algorithm (Algorithm˜1), we must verify that the maximal matrix construction in Lemma˜6 remains valid when users have multiple peak indices. Suppose user $u$ has peak indices $p,p+1,\ldots,p+m$ for some $m\geq 0$, and this information is known to the algorithm. The maximal matrix consistent with this structure should assign the same value to all peak indices, equal to the minimum of their upper confidence bounds. Formally, one could define:
$P^{t}_{u,k}=\begin{cases}\min_{i\in\{0,\ldots,m\}}\textnormal{UCB}_{u,p+i}(t),&k\in\{p,p+1,\ldots,p+m\}\\ \min_{i:k\preceq i\preceq p+m}\textnormal{UCB}_{u,i}(t),&k\prec p\\ \min_{i:p\preceq i\preceq k}\textnormal{UCB}_{u,i}(t),&p+m\prec k.\end{cases}$